Practice: Solve using dimensional analysis where appropriate. All numbers must be labeled fully.

1a) Balance: ____ C4H8 (g) + ____ O2 (g) =  ____ CO2 (g) + ____ H2O (g)


Given: If 255 g C4H8 reacts with 325 g O2, then theoretically how many grams of water vapor will form? b) Determine the limiting reactant through calculations.

c) Determine the theoretical mass (g) of water vapor that can be produced.

d) Calculate the mass (g) of the reactant in excess that will be left over when the reaction is complete.



1
Expert's answer
2021-02-09T03:45:34-0500

1 a) C4H8 + 6O2 = 4CO2 + 4H2O

255g C4H8 × "\frac{1 mol C_4H_8}{56.108gC_4H_8}" = 4.54 mol C4H8


325g O2 × "\frac{1molO_2}{32gO_2}" = 10.156 mol O2


b) There are 4.54 mol C4H8 and 10.156 mol O2

To use all the C4H8, how many moles of O2 do we need

4.54 mol C4H8 × "\frac{6molO_2}{1 mol C_4H_8}" = 27.36 mol O2

To use all the O2 , how many moles of C4H8 do we need

10.156 g O2 × "\frac{1mole C_4H_8}{6molO_2}" = 1.692 mol C4H8

We have enough C4H8 to use all the O2 but we dont have enough O2 to use all the C4H8 .

O2 is the limiting reactant .


c) How many moles of H2O can we use all the O2

10.156 mol O2 × "\frac{4molH_2O}{6mol O_2}" = 6.77 mol H2O

How many grams of H2O is this

6.77 × 18.015 = 121.96 g

121.96 g mass of the H2O we produced


d). C4H8 is the Excess reactant

How much are left over

Excess reactant = total reactant - uses reactant

Excess reactant = 4.54 - 1.692

Excess reactant = 2.848 moles of C4H8

2.848 moles C4H8 × "\frac{56.11g C_4H_8}{1moleC_4H_8}" = 159.80g

159.80 g C4H8 are left over